how to solve x+3y=11,x+4y=14
Answer by Maths68(1474)

x+4y=14 ..........(1)
x+3y=11 ..........(2)
Subtract (2) from (1)
x+4y=14
-x-3y=-11
--------------
y=3
Plug in the value of y in (1)
x+4y=14
x+4(3)=14
x+12=14
Subtract 12 from both sides
x+12-12=14-12
Cancel out 12 from Left Hand Side (LHS)
x=2
x,y = 2, 3


Check
======
Plug in the values of x and y in (1)
x+4y=14 ..........(1)
2+4(3)=14
2+12=14
14=14
Now
Plug in the values of x and y in (2)
x+3y=11 ..........(2)
2+3(3)=11
2+9=11
11=11